OpenStudy (anonymous):

solve using the miltiplication principle -1/2x=-3/4

OpenStudy (agreene):

if you mean \[\frac{-1}{2}x=\frac{-3}{4}\] then x=3/2 however, if you mean \[\frac{-1}{2x}=\frac{-3}{4}\] then x= 2/3